OpenStudy (anonymous):

assume that you are given the graph of a function explain briefly how do you know that the graph really represents a function?

OpenStudy (solomonzelman):

IS this calculus or just basic geometry?

OpenStudy (anonymous):

its for a college algebra class

OpenStudy (solomonzelman):

Lets say you have knowing these you can say that it is f(x) = (x+3)(x+1)(x-1) expand the second 2 parenthesis this is easier f(x) = (x+3)(x^2-1) f(x) = x^3+3x^2-x-3 See how I found the function using just the given graph?